The probability is

Use the normal approximation to the binomial to find the probability for n-50, p 0.6, and X = 31, Round~-value calculations to 2 decimal places and final answer to 4 decimal places. The probability is
Answer #1

\mathbf{Solution}

\mathbf{given:}p=0.6,q=0.4,n=50

np=50*0.6=30\geq 5

ng = 50 * 0.4-20 > 5

Binomial random variable is approximately normal

Mean\;(\mu)=np=30

Standard\;deviation\;(\mu)=\sqrt{npq}=\sqrt{50*0.6*0.4}=\sqrt{12}=3.4641

P(X-31) = P(30.5 〈 X < 31.5) (continuity correction)

P (30.56430 <-< 3346430)- P(0.14< z < 0.43) 3.4641 3.4641ーノ-

=P(0.43)-P(0.14)

=0.6664-0.5557

={\color{Red} 0.1107}